OpenStudy (anonymous):

Medals and a fan for help? (: 1. x+2y=23 5x+10y=55 2. 5x+4y=-83 3x-3y=-12 3. 4x+y-21 -2x+6y=9

OpenStudy (anonymous):

1.SOLUTION: x+2y=23 5x+10y=55

OpenStudy (anonymous):

2.You have to find x and y substitution is the easiest method so i will solve it by that 4y=-5x-83 y=-5/4x - 20.75 3x-3(-5/4x - 20.75)=-12 3x+3.75x+62.25=-12 6.75x=-74.25 x=-11 y=-5/4(-11) - 20.75 y=-7 Your answers are x=-11 and y=-7

OpenStudy (anonymous):

3. The solution to the problem you asked, 4x+y=21; -2+6y=9 is x = 115/24, y = 11/6.

OpenStudy (anonymous):

there u go plz fan me anf give me an medal

OpenStudy (anonymous):

I'm confused on what number one is though

OpenStudy (anonymous):

oops

OpenStudy (anonymous):